OpenStudy (anonymous):

What are the solutions of the system? y=x+5 y=x^2-x+2

OpenStudy (anonymous):

Combine the equations to get: \[x^{2} -2x -3 = y. \]

OpenStudy (anonymous):

Oops, I meant set them equal to each other.

OpenStudy (mathstudent55):

Since you have y equal to two different expressions in x, set the two expressions equal to each other. \(x + 5 = x^2 - x + 2\) Now bring all terms to the left side, and solve the quadratic equation in x. Then substitute each solution in the first original equiation to get a y value. You should end up with two ordered pairs.